OpenStudy (anonymous):

Vector Calculus: Double integral of (x/(5+xy)^2) for x=1->2, y=0->4

OpenStudy (anonymous):

\[\int\limits_{0}^{4}\int\limits_{1}^{2} \frac{ x }{ (5+xy)^2 } dydx\] I know that it's easier to handle y first, but I'm not sure how to do it.

OpenStudy (irishboy123):

yes do y first because x is a constant in that. in fact you can do each bit in your head. d/dy ( 5 + xy) ^ -1 = (-1)• ( ( 5 + xy) ^ -2 • x you see?

OpenStudy (amorfide):

integrate it treating y as a constant substitute your x values in and then integrate the new function treating x as a constant

OpenStudy (anonymous):

solve this first \[\int_{1}^{2} \frac{x}{(5+xy)^2}dy\] assuming x as a constant

OpenStudy (anonymous):

then integrate the result with respect to x now form x=0->4

OpenStudy (anonymous):

u sub? u=5+xy

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u = 5+xy du = xdy

OpenStudy (anonymous):

results in this? \[\int\limits_{0}^{4} -\frac{ 1 }{ 5+x } dx\]

OpenStudy (irishboy123):

did you apply the limits correctly? you said in the preamble: x=1->2, y=0->4

OpenStudy (anonymous):

I switched x and y